# Arcade Processes for Informed Martingale Interpolation

### v2

@@ 4.1 The one-arc FAM
We begin with showing that $(N_{t})_{t\in[T_{0},T_{1}]}$ is an $(\mathcal{F}_{t}^{I})$-martingale. We introduce the following notation: $h_{1}(t):=H_{1}^{\prime}(t)H_{2}(T_{1})-H_{1}(T_{1})H_{2}^{\prime}(t)$, $h_{2}(t):=H_{1}^{\prime}(t)H_{2}(t)-H_{1}(t)H_{2}^{\prime}(t)$, $h_{3}(t):=H_{1}(T_{1})H_{2}(t)-H_{1}(t)H_{2}(T_{1})$, and
$S(X_{0},T_{0},t):=\int_{T_{0}}^{t}\frac{(g_{0}(u)X_{0}+\mu_{A}(u))h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u.$
@@
Let $(s,t)\in[T_{0},T_{1}]^{2}$ such that $s<t$. We next show that $\mathbb{E}[N_{t}\,|\,\mathcal{F}_{s}^{I}]=N_{s}$. By the linearity property of the conditional expectation, we have
$\mathbb{E}[N_{t}\,|\,\mathcal{F}_{s}^{I}]$ $=\int_{T_{0}}^{s}\frac{I^{(1)}_{u}h_{1}(u)-M_{u}h_{2}(u)}{h_{3}(u)}\,\mathrm{d}u+\int_{s}^{t}\frac{\mathbb{E}[I^{(1)}_{u}\,|\,\mathcal{F}_{s}^{I}]h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u-M_{s}\int_{s}^{t}\frac{h_{2}(u)}{h_{3}(u)}\,\mathrm{d}u$
$\quad-(S(X_{0},T_{0},s)+S(X_{0},s,t))-g_{0}(t)X_{0}-\mu_{A}(t)+\mathbb{E}[I^{(1)}_{t}\,|\,\mathcal{F}_{s}^{I}].$ (4.45)
By Proposition 4.4, the second and the last terms in the above expression of $\mathbb{E}[N_{t}\,|\,\mathcal{F}_{s}^{I}]$ can be expressed as follows:
$\mathbb{E}[I_{t}^{(1)}\,|\,\mathcal{F}^{I}_{s}]=\left(g_{0}(t)-\frac{A_{2}(t)}{A_{2}(s)}g_{0}(s)\right)X_{0}+\left(g_{1}(t)-\frac{A_{2}(t)}{A_{2}(s)}g_{1}(s)\right)M_{s}$
$\hskip 256.0748pt+\frac{A_{2}(t)}{A_{2}(s)}I_{s}^{(1)}+\mu_{A}(t)+\frac{A_{2}(t)}{A_{2}(s)}\mu_{A}(s),$ (4.46)
$\int_{s}^{t}\frac{\mathbb{E}[I^{(1)}_{u}\,|\,\mathcal{F}_{s}^{I}]h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u=X_{0}\int_{s}^{t}\frac{g_{0}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u+M_{s}\int_{s}^{t}\frac{g_{1}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u+\int_{s}^{t}\frac{\mu_{A}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u$
$\hskip 142.26378pt+\left(I_{s}^{(1)}-g_{1}(s)M_{s}-g_{0}(s)X_{0}+\mu_{A}(s)\right)\frac{1}{A_{2}(s)}\int_{s}^{t}\frac{A_{2}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u.$ (4.47)
@@
$\mathbb{E}[I_{t}^{(1)}\,|\,\mathcal{F}^{I}_{s}]+\int_{s}^{t}\frac{\mathbb{E}[I^{(1)}_{u}\,|\,\mathcal{F}_{s}^{I}]h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u$
$=(g_{0}(t)-g_{0}(s))X_{0}+(g_{1}(t)-g_{1}(s))M_{s}+X_{0}\int_{s}^{t}\frac{g_{0}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u+M_{s}\int_{s}^{t}\frac{g_{1}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u$
$\quad\quad+\int_{s}^{t}\frac{\mu_{A}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u+I_{s}^{(1)}+\mu_{A}(s)+\mu_{A}(t)$
$=(g_{0}(t)-g_{0}(s))X_{0}+(g_{1}(t)-g_{1}(s))M_{s}+S(X_{0},s,t)+M_{s}\int_{s}^{t}\frac{g_{1}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u+I_{s}^{(1)}$
$\quad\quad+\mu_{A}(s)+\mu_{A}(t).$ (4.49)
@@
$\mathbb{E}[I_{t}^{(1)}\,|\,\mathcal{F}^{I}_{s}]+\int_{s}^{t}\frac{\mathbb{E}[I^{(1)}_{u}\,|\,\mathcal{F}_{s}^{I}]h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u$
@@
$\mathbb{E}[N_{t}\,|\,\mathcal{F}_{s}^{I}]=\int_{T_{0}}^{s}\frac{I^{(1)}_{u}h_{1}(u)-M_{u}h_{2}(u)}{h_{3}(u)}\,\mathrm{d}u-S(X_{0},T_{0},s)-g_{0}(s)X_{0}-\mu_{A}(s)+I^{(1)}_{s}=N_{s}.$ (4.52)
[…]
$[W]_{t}=\int_{T_{0}}^{t}\frac{1}{h_{2}(s)}\,\mathrm{d}[N]_{s}=\int_{T_{0}}^{t}\frac{1}{h_{2}(s)}\,\mathrm{d}[I^{(1)}]_{s}=\int_{T_{0}}^{t}\frac{h_{2}(s)}{h_{2}(s)}\,\mathrm{d}s=t-T_{0}.$ (4.53)
Since $(N_{t})$ is an $(\mathcal{F}_{t}^{I})$-martingale and $\mathbb{E}[[W]_{t}]<\infty$, $(W_{t})$ is an $(\mathcal{F}_{t}^{I})$-martingale. Hence, by Lévy’s characterization theorem, $(W_{t})_{t\in[T_{0},T_{1}]}$ is an $(\mathcal{F}_{t}^{I})$-adapted standard Brownian motion on $[T_{0},T_{1}]$. ∎
@@
$M_{t}=X_{0}+\int_{T_{0}}^{t}\frac{\sqrt{H_{1}^{\prime}(s)H_{2}(s)-H_{1}(s)H_{2}^{\prime}(s)}\mathrm{Var}[X_{1}\,|\,X_{0},I_{s}^{(1)}]}{H_{1}(T_{1})H_{2}(s)-H_{1}(s)H_{2}(T_{1})}\,\mathrm{d}W_{s}.$ (4.54)
In the case that $(I^{(1)}_{t})$ is an $X$-randomized anticipative Brownian bridge on $[T_{0},T_{1}]$, that is,
@@
where $(B_{t})_{t\geq 0}$ is a standard Brownian motion, the expressions become significantly simpler.
@@
If $(I^{(1)}_{t})_{t\in[T_{0},T_{1}]}$ is an $X$-randomized anticipative Brownian bridge on $[T_{0},T_{1}]$, the following holds:
@@
    $M_{t}=X_{0}+\int_{T_{0}}^{t}\frac{\mathrm{Var}[X_{1}\,|\,X_{0},I_{t}^{(1)}]}{T_{1}-t}\left(\frac{I_{t}^{(1)}-M_{t}}{T_{1}-t}\,\mathrm{d}t+\,\mathrm{d}I^{(1)}_{t}\right).$ (4.56)
@@
    $W_{t}:=\int_{T_{0}}^{t}\frac{I_{u}^{(1)}-M_{u}}{T_{1}-u}\,\mathrm{d}u+I^{(1)}_{t}-X_{0},$ (4.57)
    is an $(\mathcal{F}_{t}^{I})$-adapted standard Brownian motion on $[T_{0},T_{1}]$.
@@
It suffice to set $H_{1}(x)=x$, $H_{2}(x)=1$, $g_{0}(t)=(T_{1}-t)/(T_{1}-T_{0})$, and $g_{1}(t)=(t-T_{0})/(T_{1}-T_{0})$ in Eqs. (4.54) and (4.43). ∎
In case that $X_{0}=0$, $X_{1}$ is centered around 00, and $T_{0}=0$, the randomized anticipative Brownian bridge gives rise to the martingale developed in [7], i.e,
$M_{t}=\int_{0}^{t}\frac{\mathrm{Var}[X_{1}\,|\,I_{s}^{(1)}]}{T_{1}-s}\,\mathrm{d}W_{s}.$ (4.58)
@@
Example 4.16.
Let $D_{t}=tB_{t}$ where $(B_{t})_{t\geqslant 0}$ is a standard Brownian motion. Then, $K_{D}(x,y)=\min(x,y)^{2}\max(x,y)$, and so $H_{1}(x)=x^{2}$, $H_{2}(x)=x$. The standard AP on $[T_{0},T_{1}]$ driven by $(D_{t})$ is given by
$A_{t}^{(1)}=D_{t}-\frac{T_{1}t-t^{2}}{T_{1}T_{0}-T_{0}^{2}}D_{T_{0}}-\frac{t^{2}-tT_{0}}{T_{1}^{2}-T_{1}T_{0}}D_{T_{1}}.$ (4.59)
We have
$K_{A}(x,y)=\frac{\min(x,y)(\min(x,y)-T_{0})\max(x,y)(T_{1}-\max(x,y))}{T_{1}-T_{0}},$ (4.60)
$I_{t}^{(1)}=D_{t}-\frac{T_{1}t-t^{2}}{T_{1}T_{0}-T_{0}^{2}}(D_{T_{0}}-X_{0})-\frac{t^{2}-tT_{0}}{T_{1}^{2}-T_{1}T_{0}}(D_{T_{1}}-X_{1}).$ (4.61)
$M_{t}=X_{0}+\int_{T_{0}}^{t}\frac{\mathrm{Var}[X_{1}\,|\,X_{0},I_{t}^{(1)}]s}{T_{1}^{2}s-s^{2}T_{1}}\,\mathrm{d}W_{s}=X_{0}+\int_{T_{0}}^{t}\frac{\mathrm{Var}[X_{1}\,|\,X_{0},I_{s}^{(1)}]}{T_{1}^{2}-sT_{1}}\,\mathrm{d}W_{s}.$ (4.62)
Example 4.17.
Let $(D_{t})$ be the Ornstein-Uhlenbeck process satisfying the SDE
$D_{t}=\int_{0}^{t}\theta\left(\mu-D_{s}\right)\,\mathrm{d}s+\int_{0}^{t}\sigma\,\mathrm{d}B_{s},$ (4.63)
$A_{t}^{(1)}=D_{t}-\frac{\mathrm{e}^{\theta(T_{1}-t)}-\mathrm{e}^{-\theta(T_{1}-t)}}{\mathrm{e}^{\theta(T_{1}-T_{0})}-\mathrm{e}^{-\theta(T_{1}-T_{0})}}D_{T_{0}}-\frac{\mathrm{e}^{\theta(t-T_{0})}-\mathrm{e}^{-\theta(t-T_{0})}}{\mathrm{e}^{\theta(T_{1}-T_{0})}-\mathrm{e}^{-\theta(T_{1}-T_{0})}}D_{T_{1}}.$ (4.64)
$I_{t}^{(1)}=D_{t}-\frac{\mathrm{e}^{\theta(T_{1}-t)}-\mathrm{e}^{-\theta(T_{1}-t)}}{\mathrm{e}^{\theta(T_{1}-T_{0})}-\mathrm{e}^{-\theta(T_{1}-T_{0})}}(D_{T_{0}}-X_{0})-\frac{\mathrm{e}^{\theta(t-T_{0})}-\mathrm{e}^{-\theta(t-T_{0})}}{\mathrm{e}^{\theta(T_{1}-T_{0})}-\mathrm{e}^{-\theta(T_{1}-T_{0})}}(D_{T_{1}}-X_{1}).$ (4.65)
$M_{t}=X_{0}+\int_{T_{0}}^{t}\frac{\sigma\mathrm{Var}[X_{1}\,|\,X_{0},I_{t}^{(1)}]}{\frac{\sigma^{2}}{2\theta}\left(\mathrm{e}^{\theta(T_{1}-s)}-\mathrm{e}^{\theta(s-T_{1})}\right)}\,\mathrm{d}W_{s}=X_{0}+\frac{2\theta}{\sigma}\int_{T_{0}}^{t}\frac{\mathrm{Var}[X_{1}\,|\,X_{0},I_{t}^{(1)}]}{\mathrm{e}^{\theta(T_{1}-s)}-\mathrm{e}^{\theta(s-T_{1})}}\,\mathrm{d}W_{s}.$ (4.66)
In general, one can expect that for a given coupling $\pi^{X}$, $M_{t}=\mathbb{E}[X_{1}\,|\,\mathcal{F}_{t}^{I}]$ has no explicit analytical expression, even in the case where $M_{t}=\mathbb{E}[X_{1}\,|\,X_{0},I_{t}^{(1)}]$. Nonetheless, we give an example where $M_{t}=\mathbb{E}[X_{1}\,|\,X_{0},I_{t}^{(1)}]$ and its SDE are explicit.
@@
$X_{1}\,|\,X_{0}=\left\{\begin{array}[]{ll}X_{0}+1&\mbox{with probability }\frac{1}{2},\\ \\ X_{0}-1&\mbox{with probability }\frac{1}{2}.\end{array}\right.$ (4.67)
Clearly, $\mathbb{E}[X_{1}\,|\,X_{0}]=X_{0}$. Let $(B_{t})_{t\geqslant 0}$ be a standard Brownian motion and
@@
Moreover, let $\phi$ denote the density function of the measure $\mathcal{N}\left(0,\frac{(T_{1}-t)(t-T_{1})}{T_{1}-T_{0}}\right)$, and recall that $g_{1}(x)=(t-T_{0})/(T_{1}-T_{0})$. By Proposition 4.5, we have:
@@
[…]
Figure 12: Paths of the $X$-RAP $(I_{t}^{(1)})$, c.f., Example 4.18.
Figure 13: Paths of the associated FAM $(M_{t})$, c.f., Example 4.18.

### v3

@@ 5.1 The one-arc FAM
We begin with showing that $(N_{t})_{t\in[T_{0},T_{1}]}$ is an $(\mathcal{F}_{t}^{I})$-martingale. We introduce the following notation: $h_{1}(t)\mathrel{\mathop{\ordinarycolon}}=H_{1}^{\prime}(t)H_{2}(T_{1})-H_{1}(T_{1})H_{2}^{\prime}(t)$, $h_{2}(t)\mathrel{\mathop{\ordinarycolon}}=H_{1}^{\prime}(t)H_{2}(t)-H_{1}(t)H_{2}^{\prime}(t)$, $h_{3}(t)\mathrel{\mathop{\ordinarycolon}}=H_{1}(T_{1})H_{2}(t)-H_{1}(t)H_{2}(T_{1})$, and
$S(X_{0},T_{0},t)\mathrel{\mathop{\ordinarycolon}}=\int_{T_{0}}^{t}\frac{(g_{0}(u)X_{0}+\mu_{A}(u))h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u.$
@@
Let $(s,t)\in[T_{0},T_{1}]^{2}$ such that $s<t$. We next show that $\mathbb{E}[N_{t}\mid\mathcal{F}_{s}^{I}]=N_{s}$. By the linearity property of the conditional expectation, we have
$\mathbb{E}[N_{t}\mid\mathcal{F}_{s}^{I}]=\int_{T_{0}}^{s}\frac{I^{(1)}_{u}h_{1}(u)-M_{u}h_{2}(u)}{h_{3}(u)}\,\mathrm{d}u+\int_{s}^{t}\frac{\mathbb{E}[I^{(1)}_{u}\mid\mathcal{F}_{s}^{I}]h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u$
$\quad-M_{s}\int_{s}^{t}\frac{h_{2}(u)}{h_{3}(u)}\,\mathrm{d}u-(S(X_{0},T_{0},s)+S(X_{0},s,t))-g_{0}(t)X_{0}-\mu_{A}(t)+\mathbb{E}[I^{(1)}_{t}\mid\mathcal{F}_{s}^{I}].$ (5.44)
By Proposition 5.30, the second and the last terms in the above expression of $\mathbb{E}[N_{t}\mid\mathcal{F}_{s}^{I}]$ can be expressed as follows:
$\mathbb{E}[I_{t}^{(1)}\mid\mathcal{F}^{I}_{s}]=\left(g_{0}(t)-\frac{A_{2}(t)}{A_{2}(s)}g_{0}(s)\right)X_{0}+\left(g_{1}(t)-\frac{A_{2}(t)}{A_{2}(s)}g_{1}(s)\right)M_{s}$
$\hskip 227.62204pt+\frac{A_{2}(t)}{A_{2}(s)}I_{s}^{(1)}+\mu_{A}(t)+\frac{A_{2}(t)}{A_{2}(s)}\mu_{A}(s),$ (5.45)
$\int_{s}^{t}\frac{\mathbb{E}[I^{(1)}_{u}\mid\mathcal{F}_{s}^{I}]h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u=X_{0}\int_{s}^{t}\frac{g_{0}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u+M_{s}\int_{s}^{t}\frac{g_{1}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u$
$+\int_{s}^{t}\frac{\mu_{A}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u+\frac{\left(I_{s}^{(1)}-g_{1}(s)M_{s}-g_{0}(s)X_{0}+\mu_{A}(s)\right)}{A_{2}(s)}\int_{s}^{t}\frac{A_{2}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u.$ (5.46)
@@
$\mathbb{E}[I_{t}^{(1)}\mid\mathcal{F}^{I}_{s}]+\int_{s}^{t}\frac{\mathbb{E}[I^{(1)}_{u}\mid\mathcal{F}_{s}^{I}]h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u$
$=(g_{0}(t)-g_{0}(s))X_{0}+(g_{1}(t)-g_{1}(s))M_{s}+X_{0}\int_{s}^{t}\frac{g_{0}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u$
$\hskip 65.44142pt+M_{s}\int_{s}^{t}\frac{g_{1}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u+\int_{s}^{t}\frac{\mu_{A}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u+I_{s}^{(1)}+\mu_{A}(s)+\mu_{A}(t)$
$=(g_{0}(t)-g_{0}(s))X_{0}+(g_{1}(t)-g_{1}(s))M_{s}+S(X_{0},s,t)$
$\hskip 170.71652pt+M_{s}\int_{s}^{t}\frac{g_{1}(u)h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u+I_{s}^{(1)}+\mu_{A}(s)+\mu_{A}(t).$ (5.48)
@@
$\mathbb{E}[I_{t}^{(1)}\mid\mathcal{F}^{I}_{s}]+\int_{s}^{t}\frac{\mathbb{E}[I^{(1)}_{u}\mid\mathcal{F}_{s}^{I}]h_{1}(u)}{h_{3}(u)}\,\mathrm{d}u$
@@
$\mathbb{E}[N_{t}\mid\mathcal{F}_{s}^{I}]=\int_{T_{0}}^{s}\frac{I^{(1)}_{u}h_{1}(u)-M_{u}h_{2}(u)}{h_{3}(u)}\,\mathrm{d}u-S(X_{0},T_{0},s)-g_{0}(s)X_{0}-\mu_{A}(s)+I^{(1)}_{s}=N_{s}.$ (5.51)
Moreover, $\mathbb{E}[\,\mathinner{\!\left\lvert N_{t}\right\rvert}$$\infty$ for all $t\in[T_{0},T_{1}]$. Hence $(N_{t})_{t\in[T_{0},T_{1}]}$ is an $(\mathcal{F}_{t}^{I})$-martingale. We now show that $(W_{t})$ is an $(\mathcal{F}_{t}^{I})$-adapted standard Brownian motion on $T_{0},T_{1}]$. We compute the quadratic variation of $(W_{t})$ and obtain
$W]_{t}=\int_{T_{0}}^{t}\frac{1}{h_{2}(s)}\,\mathrm{d}[N]_{s}=\int_{T_{0}}^{t}\frac{1}{h_{2}(s)}\,\mathrm{d}[I^{(1)}]_{s}=\int_{T_{0}}^{t}\frac{h_{2}(s)}{h_{2}(s)}\,\mathrm{d}s=t-T_{0}.$ (5.52)
Since $(N_{t})$ is an $(\mathcal{F}_{t}^{I})$-martingale and $\mathbb{E}[[W]_{t}]<\infty$, $(W_{t})$ is an $(\mathcal{F}_{t}^{I})$-martingale. Hence, by Lévy’s characterisation theorem (Theorem 7.1 in [27]), $(W_{t})_{t\in[T_{0},T_{1}]}$ is an $(\mathcal{F}_{t}^{I})$-adapted standard Brownian motion on $T_{0},T_{1}]$. ∎
@@
$M_{t}=X_{0}+\int_{T_{0}}^{t}\frac{\mathrm{Var}[X_{1}\mid X_{0},I_{s}^{(1)}]\sqrt{H_{1}^{\prime}(s)H_{2}(s)-H_{1}(s)H_{2}^{\prime}(s)}}{H_{1}(T_{1})H_{2}(s)-H_{1}(s)H_{2}(T_{1})}\,\mathrm{d}W_{s}.$ (5.53)
In the case that $(I^{(1)}_{t})$ is an $X$-randomised anticipative Brownian bridge on $T_{0},T_{1}]$, that is,
@@
where $(B_{t})_{t\geqslant 0}$ is a standard Brownian motion, the expressions become significantly simpler.
@@
If $(I^{(1)}_{t})_{t\in[T_{0},T_{1}]}$ is an $X$-randomised anticipative Brownian bridge on $T_{0},T_{1}]$, the following holds:
@@
    $M_{t}=X_{0}+\int_{T_{0}}^{t}\frac{\mathrm{Var}[X_{1}\mid X_{0},I_{t}^{(1)}]}{T_{1}-t}\left(\frac{I_{t}^{(1)}-M_{t}}{T_{1}-t}\,\mathrm{d}t+\,\mathrm{d}I^{(1)}_{t}\right)$. (5.55)
@@
    $W_{t}\mathrel{\mathop{\ordinarycolon}}=\int_{T_{0}}^{t}\frac{I_{u}^{(1)}-M_{u}}{T_{1}-u}\,\mathrm{d}u+I^{(1)}_{t}-X_{0},$ (5.56)
    is an $(\mathcal{F}_{t}^{I})$-adapted standard Brownian motion on $T_{0},T_{1}]$.
@@
It suffices to set $H_{1}(x)=x$, $H_{2}(x)=1$, $g_{0}(t)=(T_{1}-t)/(T_{1}-T_{0})$, and $g_{1}(t)=(t-T_{0})/(T_{1}-T_{0})$ in Eqs. (5.53) and (5.42). ∎
When $X_{0}=0$, $X_{1}$ is centred around 0, and $T_{0}=0$, the randomised anticipative Brownian bridge gives rise to the martingale developed in [7], i.e.,
$M_{t}=\int_{0}^{t}\frac{\mathrm{Var}[X_{1}\mid I_{s}^{(1)}]}{T_{1}-s}\,\mathrm{d}W_{s}.$ (5.57)
@@
In general, one can expect that for a given coupling $\pi^{X}$, $M_{t}=\mathbb{E}[X_{1}\mid\mathcal{F}_{t}^{I}]$ has no explicit analytical expression, even in the case where $M_{t}=\mathbb{E}[X_{1}\mid X_{0},I_{t}^{(1)}]$. Nonetheless, we give an example where $M_{t}=\mathbb{E}[X_{1}\mid X_{0},I_{t}^{(1)}]$ and its SDE are explicit.
@@
$X_{1}\mid X_{0}=\left\{\begin{array}[]{ll}X_{0}+1&\mbox{with probability }\frac{1}{2},\\ \\ X_{0}-1&\mbox{with probability }\frac{1}{2}.\end{array}\right$. (5.58)
Clearly, $\mathbb{E}[X_{1}\mid X_{0}]=X_{0}$. Let $(B_{t})_{t\geqslant 0}$ be a standard Brownian motion and
@@
Moreover, let $\phi$ denote the density function of the Gaussian measure
$\mathcal{N}\left(0,(T_{1}-t)(t-T_{1})/(T_{1}-T_{0})\right)$, and recall that $g_{1}(x)=(t-T_{0})/(T_{1}-T_{0})$. By Proposition 5.31, we have:
@@
$M_{t}=X_{0}+\mathop{\text{\LARGE$\int_{\text{\normalsize$\scriptstyle\kern-2.04861ptT_{0}$}}^{\text{\normalsize$\scriptstyle t$}}$}}\nolimits\frac{\operatorname{sech}^{2}\left(\frac{I_{s}-X_{0}}{T_{1}-s}\right)}{T_{1}-s}\,\mathrm{d}W_{s}.$ (5.63)
Figure 5: Paths of the $X$-RAP $(I_{t}^{(1)})$, cf., Example 5.40.
Figure 6: Paths of the associated FAM $(M_{t})$, cf., Example 5.40.